Use the given information to answer parts (a) through (d) Sample statistics x -357,s, -39.n, -17 345 25. n 8 Assume ? ? to-1.714, to 1.714 (a) Find the test statistic
Answer #1

Here pooled standard deviation

sp = sqrt [{(n1 -1)s12 + (n2 -1)s22}/ (n1 + n2 - 2)]

sp = sqrt [(16 *3 .9 * 3.9 + 7 * 2.5 * 2.5)/ (17 + 8 -2)] = 3.5331

Here test statistic

t = (x?1 - x?2)/ [sp * sqrt(1/n1 + 1/n2)]

t = (35.7 - 34.5)/ [3.5331 * sqrt (1/17 + 1/9)]

t = 1.2/ 1.4565

t = 0.8239

Here t < t0 so we will fail to rejecthe null hypothesis and say both means are equal here.